OpenStudy (anonymous):

Can somebody help me please Y = -x + 150 Y = -19/30x + 106

OpenStudy (anonymous):

y + x + 150 = 0 Y + 19/30x + 106 = 0 Now subtract both the equations

OpenStudy (anonymous):

11/30x + 44 = 0 11/30x = -44 11/-44 = 30x Now u can get x value and substitute that in any of the equation to get y value

OpenStudy (anonymous):

Can you show me : ) @Gatorgirl

OpenStudy (gatorgirl):

I wish, But I suck at Math. I'm sorry. But It seems like @ganibl got it. :)

OpenStudy (anonymous):

-1/4 = 30x = > x = -1/120 Substitute this in the equations given (any one is ok) y = -x + 150 = 1/120 + 150 = 18120/120 = 151

OpenStudy (anonymous):

Is this a answer ( 120,30)

OpenStudy (anonymous):

@ganibl

OpenStudy (anonymous):

yes perfect :)

OpenStudy (anonymous):

I am not sure what the question is asked for so I did elimination method :) But if u have options u can directly check by substituting them in the equation and see if it holds true for those values :)

OpenStudy (anonymous):

What did you get @ganibl

OpenStudy (anonymous):

(-1/120 , 151) But I might be wrong some where